# thanks: These authors contributed equally

### v1

@@ Abstract
Graphs are used to represent and analyze data in domains as diverse as physics, biology, chemistry, planetary science, and the social sciences. Across domains, random graph models relate generative processes to expected graph properties, and allow for sampling from distinct ensembles. Here we introduce a new random graph model, inspired by assembly theory, and characterize the graphs it generates. We show that graphs generated using our method represent a diverse ensemble, characterized by a broad range of summary statistics, unexpected even in graphs with identical degree sequences. Finally we demonstrate that the distinct properties of these graphs are enabled by historical contingencies during the generative process. These results lay the foundation for further development of novel sampling methods based on assembly theory with applications to drug discovery and materials science.
@@
Here we introduce a new generative procedure for sampling graphs, based on assembly theory, and characterize the ensemble it produces. Assembly theory (AT) is a new theoretical framework for characterizing selection across diverse objects, most importantly in molecules [15]. In AT, objects are defined as finite in extent, persistent in time, distinguishable, and decomposable into basic building blocks. Graphs satisfy all of these conditions (with the possible exception of persistence in time, as the ontological status of mathematical entities is debated). A central quantity in AT is the assembly index of objects, which is the minimum number of joining operations required to construct the object from basic components, in which recursively generated objects can be reused as a single joining operation [15, 16]. The assembly index can be measured empirically for molecules, enabling novel approaches to quantify life detection, evolutionary relationships, and material characteristics [17, 18, 19, 20]. The definition of the assembly index implies a constructive procedure that when paired with empirical data can produce structures with desirable properties, such as similarity to target compounds or increased “drug-likeness” [21, 22]. Here we generalize this approach to graphs by defining the constructive steps, and then characterize the ensemble of graphs induced by this procedure. This procedure is distinctive from other random graph models because it both recursively reuses substructures (without preserving the entire adjacency matrix) and is dynamic in nature [13, 12]. Here we describe the sampling procedure in detail, provide an algorithmic implementation, and show that the graphs generated by this procedure are exceptional (compared to graphs with identical degree sequences) based on their global topological properties such as mean betweenness, clustering coefficient, and algebraic connectivity [23, 11]. Finally we demonstrate how the contingency induced by this procedure enables the generation of diverse samples with these exceptional properties, and discuss the implications to exploration of chemical space.
@@
Model Description—The random graph assembly process begins with a multiset $\mathbb{G}$ containing simple graphs. At each iteration, two graphs are selected from $\mathbb{G}$, with replacement. The first graph, $L$, is chosen through a biased selection process: with probability $p$, $L$ is selected uniformly at random from $\mathbb{G}_{\max}$ (a subset of $\mathbb{G}$ containing only graphs with the most nodes); with probability $1-p$, $L$ is selected uniformly at random from $\mathbb{G}$. The second graph, $R$, is always selected uniformly at random from $\mathbb{G}$. It is possible for $L$ and $R$ to be identical graphs. The process then attempts to merge $M$ pairs of vertices between $L$ and $R$, with a vertex pair always containing one node in $L$ and one node in $R$. No vertex may be in more than one pair. When two vertices are merged, the resulting vertex inherits all edges from both original vertices, while maintaining the properties of a simple graph: parallel edges are combined, and self-loops are prohibited. The resulting graph is then added to $\mathbb{G}$. This entire process is repeated for $N$ iterations. At the end of the run $\mathbb{G}$ contains a variety of randomly generated graphs. We focus our analysis on the largest graph in the set at each iteration, though the properties of the entire set are an interesting topic for future study. Fig. 1 illustrates this procedure conceptually (panel a), and gives an example trajectory of largest graphs assembled for 10 iterations of this process (panel b).
@@
Specifying an instance of this algorithm requires specifying an initial multiset of graphs $\mathbb{G}$, the number of iterations $N$, the probability $p$ of selecting $L$ from $\mathbb{G}_{\max}$, and a procedure for selecting $M$. For the results presented here we always initialize $\mathbb{G}$ with a path graph of two nodes and a path graph of three nodes (see Appendix A). We chose $M$ at each iteration uniformly at random from the discrete range $[1,m]$, where $M_{\max}=\min(|V(L)|,|V(R)|)$. Higher values of $p$ yield larger graphs in fewer iterations but reduce the diversity of accessible graphs.
@@
We repeated this analysis for 1000 randomly assembled graphs, calculating these parameters after 25 iterations, using 1000 controls for each graph. We found that the assembled graphs frequently exhibited values in the most extreme upper/lower tails of these parameters as compared to randomized configuration models with the same degree sequences. To quantify this we calculated the $Z$ score of the relevant statistic for randomly assembled graph compared in the ensemble of the random configuration graphs derived from it. The results are shown in Fig. 2 ($p=0.5$, and $M\in[1,3]$). These results demonstrate that our algorithm can sample graphs which include typical random graphs but often represent highly selected graphs, indicating the algorithm is consistently sampling graphs that represent extreme cases compared to other sampling procedures.
A key feature of this algorithm is the contingency that is induced within a single trajectory. To explore how this feature controls the properties of the produced graphs we performed simulations to resample from trajectories with some of this contingency removed in two different ways. First we explored the effect of resampling the trajectories to understand how the historical features controlled the entire ensemble of possible trajectories. Specifically we do the following: for a single trajectory of the algorithm (a reference trajectory), we generate a new input multiset, $\mathbb{R}^{\prime}_{c}$, which is the final multiset of graphs, $\mathbb{R}$, accumulated in the reference trajectory with the graphs accumulated in the last $c$ steps removed. We then run the algorithm with $\mathbb{R}^{\prime}_{c}$ for $c$ steps, yielding a new trajectory with an identical history for the first $N-c$ steps, but allowing for a distinct final $c$ steps. We refer to this trajectory as having a history loss of $c$ steps. Fig. 3 shows the mean degree of 1000 trajectories for different values of $c$ ($p$ = 0.5, $M\in[1:2]$). Larger values of $c$ enable larger variations in the final properties of the graph, consistent with the idea that fluctuations in the trajectory enable sampling of distinct sections of the graph space later in the trajectory.
@@
Next we explored the effect of removing the influence of previous graphs on the produced graph—effectively erasing history within a trajectory—while controlling for the sequential construction process (including the controlling for the mean degree). Specifically we generated a reference trajectory, tracking the size (number of nodes and edges) of each pair of graphs that was joined. Then we run a variation of the algorithm where after $L$ $R$ are chosen their connectivity is randomized, such that the number of nodes and edges are preserved and the graph is still connected, but otherwise totally random (see Appendix A). Fig. 4 shows the difference between our algorithm and this random control for the global measures we calculated. The distributions of graph properties for the original and randomized methods are all statistically significant.
Figure 4: Topological properties of randomly assembled graphs depend on history and dynamics. The distribution of graph topological properties is shown for each property for the randomly assembled graphs, and for the control that randomizes connectivity (preserving number of nodes and edges) at each step after selecting graphs to merge. This controls for the size of the joined graphs and the sequence of joining operations. The distributions for the original randomly assembled graphs and control graphs are distinct, indicating that the historical dependence of graph reuse is partially responsible for the observed topological properties.
@@ Appendix A Graph Assembly Algorithm Details
Our graph assembly requires combining graphs with at least 1 edge to allow the possibility of the resulting graph to be larger than the two merged graphs (e.g., consider the scenario with a singleton node graph $L_{\text{singleton}}$. Because merging two graphs requires $M\geq 1$, joining $L_{\text{singleton}}$ with any graph $R$ will always result in $R$). While the simplest such set of initial graphs would be two path graphs of two nodes each, this will always lead to a path graph of three nodes before complexifying further, hence we initialize our algorithm with two path graphs, of three nodes and two nodes respectively.
@@
When multiple graphs in $\mathbb{G}$ meet the condition for graph with the largest graph (by number of nodes), one is chosen at random.

### v2

@@ Abstract
Graphs are used to represent and analyze data in physics, biology, chemistry, planetary science, and the social sciences. Across domains, random graph models relate generative processes to expected graph properties, and allow for sampling from distinct ensembles. Here we introduce a new random graph model, inspired by assembly theory, and characterize the graphs it produces. We show that graphs generated using our method represent a diverse ensemble, characterized by a broad range of summary statistics, unexpected even in graphs with identical degree sequences. Finally we demonstrate that the distinct properties of these graphs are enabled by historical contingencies during the generative process. These results lay the foundation for further development of novel sampling methods based on assembly theory with applications to drug discovery and materials science.
@@
Here we introduce a new generative procedure for sampling graphs, based on assembly theory, and characterize the ensemble it produces. Assembly theory (AT) is a new theoretical framework for characterizing selection across diverse objects, most importantly in molecules [15]. In AT, objects are defined as finite in extent, persistent in time, distinguishable, and decomposable into basic building blocks. Graphs satisfy all of these conditions (with the possible exception of persistence in time, as the ontological status of mathematical entities is debated). A central quantity in AT is the assembly index of objects, which is the minimum number of joining operations required to construct the object from basic components, in which recursively generated objects can be reused as a single joining operation [15, 16]. The assembly index can be measured empirically for molecules, enabling novel approaches to quantify life detection, evolutionary relationships, and material characteristics [17, 18, 19, 20]. The definition of the assembly index implies a constructive procedure that when paired with empirical data can produce structures with desirable properties, such as similarity to target compounds or increased “drug-likeness” [21, 22]. Here we generalize this approach to graphs by defining the constructive steps, and then characterize the ensemble of graphs induced by this procedure. This procedure is distinctive from other random graph models because it both recursively reuses substructures (without preserving the entire adjacency matrix) and is dynamic in nature [13, 12]. Here we describe the sampling procedure in detail, provide an algorithmic implementation [23], and show that the graphs generated by this procedure are exceptional (compared to graphs with identical degree sequences) based on their global topological properties such as mean betweenness, clustering coefficient, and algebraic connectivity [24, 11]. Finally we demonstrate how the contingency induced by this procedure enables the generation of diverse samples with these exceptional properties, and discuss the implications to exploration of chemical space.
@@
Model Description—The random graph assembly process begins with a multiset $\mathbb{G}$ containing simple graphs. At each iteration, two graphs are selected from $\mathbb{G}$, with replacement. The first graph, $L$, is chosen through a biased selection process: with probability $p$, $L$ is selected uniformly at random from $\mathbb{G}_{\max}$ (a subset of $\mathbb{G}$ containing only graphs with the most nodes); with probability 1p1-p, $L$ is selected uniformly at random from $\mathbb{G}$. The second graph, $R$, is always selected uniformly at random from $\mathbb{G}$. It is possible for $L$ and $R$ to be identical graphs. The process then attempts to merge $M$ pairs of vertices between $L$ and $R$, with a vertex pair always containing one node in $L$ and one node in $R$. No vertex may be in more than one pair. When two vertices are merged, the resulting vertex inherits all edges from both original vertices, while maintaining the properties of a simple graph: parallel edges are combined, and self-loops are prohibited. The resulting graph is then added to $\mathbb{G}$. This entire process is repeated for $N$ iterations. At the end of the run $\mathbb{G}$ contains a variety of randomly generated graphs. We focus our analysis on the largest graph in the set at each iteration, though the properties of the entire set are an interesting topic for future study. Fig. 1 illustrates this procedure conceptually (panel a), and gives an example trajectory of largest graphs assembled for 10 iterations of this process (panel b).
@@
Specifying an instance of this algorithm requires specifying an initial multiset of graphs $\mathbb{G}$, the number of iterations $N$, the probability $p$ of selecting $L$ from $\mathbb{G}_{\max}$, and a procedure for selecting $M$. For the results presented here we always initialize $\mathbb{G}$ with a path graph of two nodes and a path graph of three nodes (see Appendix A). We chose $M$ at each iteration uniformly at random from the discrete range $[1,m]$. Higher values of $p$ yield larger graphs in fewer iterations but reduce the diversity of accessible graphs.
@@
We repeated this analysis for 1000 randomly assembled graphs, calculating these parameters after 25 iterations, using 1000 controls for each graph. We found that the assembled graphs frequently exhibited values in the most extreme upper/lower tails of these parameters as compared to randomized configuration models with the same degree sequences. To quantify this we calculated the $Z$ score of the relevant statistic for randomly assembled graph compared in the ensemble of the random configuration graphs derived from it. The results are shown in Fig. 2 ($p=0.5$, and $M\in[1,3]$). These results demonstrate that our algorithm can sample graphs which include typical random graphs but often represent extremely atypical graphs, indicating the algorithm is consistently sampling graphs that represent extreme cases compared to other sampling procedures.
@@
A key feature of this algorithm is the contingency that is induced within a single trajectory. To explore how this feature controls the properties of the produced graphs we performed simulations to resample from trajectories with some of this contingency removed in two different ways. First we explored the effect of resampling the trajectories to understand how the historical features controlled the entire ensemble of possible trajectories. Specifically we do the following: for a single trajectory of the algorithm (a reference trajectory), we generate a new input multiset, $\mathbb{R}^{\prime}_{c}$, which is the final multiset of graphs, $\mathbb{R}$, accumulated in the reference trajectory with the graphs accumulated in the last $c$ steps removed. We then run the algorithm with $\mathbb{R}^{\prime}_{c}$ for $c$ steps, yielding a new trajectory with an identical history for the first NcN-c steps, but allowing for a distinct final $c$ steps. We refer to this trajectory as having a history loss of $c$ steps. Fig. 3 shows the distribution of mean degrees for 1000 trajectories for different values of $c$ ($p$ = 0.5, $M\in[1:2]$). The intervals indicate the inter-quantile range between 1% and 99% for the distributions. Larger values of $c$ enable larger variations in the final properties of the graph, consistent with the idea that fluctuations in the trajectory enable sampling of distinct sections of the graph space later in the trajectory.
Figure 4: Topological properties of randomly assembled graphs depend on history and dynamics. The distribution of graph topological properties is shown for each property for the randomly assembled graphs, and for the control that randomizes connectivity (preserving number of nodes and edges) at each step after selecting graphs to merge. This controls for the size of the joined graphs and the sequence of joining operations. The distributions for the original randomly assembled graphs (shown in opaque colors) and control graphs (outlined in black) are distinct, indicating that the historical dependence of graph reuse is partially responsible for the observed topological properties.
Next we explored the effect of removing the influence of previous graphs on the produced graph—effectively erasing history within a trajectory—while controlling for the sequential construction process (including the controlling for the mean degree). Specifically we generated a reference trajectory, tracking the size (number of nodes and edges) of each pair of graphs that was joined. Then we run a variation of the algorithm where after $L$ & $R$ are chosen their connectivity is randomized, such that the number of nodes and edges are preserved and the graph is still connected, but otherwise random (see Appendix A). Fig. 4 shows the difference between our algorithm and this random control for the global measures we calculated with the original graphs in solid colors, and the randomized controls shown in black. The distributions of graph properties for the original and randomized methods are all statistically significant.
@@ Acknowledgements.
Code Availability. Code associated with this manuscript is available at https://github.com/mathis-group/AssemblingGraphs.jl
[…]
@@ Appendix A Graph Assembly Algorithm Details
Our graph assembly requires combining graphs with at least 1 edge to allow the possibility of the resulting graph to be larger than the two merged graphs (e.g., consider the scenario with a singleton node graph $L_{\text{singleton}}$. Because merging two graphs requires $M\geq 1$, joining $L_{\text{singleton}}$ with any graph $R$ will always result in $R$). While the simplest such set of initial graphs would be two path graphs of two nodes each, this will always lead to a path graph of three nodes before complexifying further, hence we initialize our algorithm with two path graphs, of three nodes and two nodes respectively. We anticipate future work could consider optimization of this initial set to sample different ensembles of assembled graphs
@@
When multiple graphs in $\mathbb{G}$ meet the condition for graph with the largest graph (by number of nodes), one is chosen uniformly at random.